# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

### v3

@@ 5 Main Results
[…]
@@
Our human evaluation results in table 3 reveal three key findings. First, ChatGPT and ChatGPT-Instruct achieve similar performance in filtering explicit personae and ranking opinions, while Mistral achieves lower results. While ChatGPT excels slightly in ranking, ChatGPT-Instruct performs slightly better in explicit personae selection. Three models proficiently filter unnecessary explicit personae, but ranking opinions poses a more challenging task intuitively and empirically, with a common error being the inconsistent relevance ranking of opinions, sometimes misplacing high-level relevance. Second, four models effectively generate reasonable thoughts leading to the final answer, and GPT-4 performs the best. Finally, ChatGPT and ChatGPT-Instruct follow our instructions to explain and analyze the explicit and implicit personae provided one by one with CoO significantly better than GPT-4 and Mistral, achieving nearly perfect scores of $3$. We hypothesize that this is because ChatGPT and ChatGPT-Instruct excel in following instructions, while GPT-4 is optimized for completing texts.
@@ 6 Discussion
We discuss the main analyses in this section. Extra analyses are presented in appendix B.
6.1 Methodology Analysis
Ablation of FEA.
@@ 6.1 Methodology Analysis
To gauge the impact of filtering unnecessary explicit personae (FEA) on performance, we experiment with applying FEA exclusively to the baseline DIO-top8 Hwang et al. (2023), denoted as DIO-top8 + FEA in table 1. The results indicate enhancements with DIO-top8 + FEA achieving a 0.8%, 1.3%, 1.9% accuracy performance boost on ChatGPT, ChatGPT-Instruct, and Mistral respectively. This underscores the effectiveness of eliminating irrelevant explicit personae in improving the models’ ability to understand and predict human opinions.
FEA via Topics.
@@
To understand the explicit personae filtered by LLMs across various topics, we document the top-3 removed personae in section B.5. We observe that "Citizenship" is consistently the most frequently removed attribute, followed by "Race". This could be due to LLMs treating these as sensitive information, prioritizing respect and unbiased text generation. Another explanation may be the lack of correlation between citizenship/race and opinions in the US-centric OpinionQA dataset. Additionally, we also see that ChatGPT often categorizes ‘‘Marital status" as non-useful, ChatGPT-Instruct commonly removes ‘‘Frequency of religious attendance", and ‘‘Gender" got removed by Mistral, revealing potential biases in LLMs.
LLMtop-$K$ versus Top-$K$.
@@
From table 1, DIO-LLMtop8 outperforms DIO-top8 by $1.6\%$, $1.6\%$, $3.8\%$ accuracy on ChatGPT, ChatGPT-Instruct, Mistral confirming that prioritizing meaning and usefulness improves opinion prediction. One possible explanation for this can be the orders ranked by semantic similarity scores only consider ranking with the input questions Hwang et al. (2023), while our orders consider both input questions and their answer choices (fig. 1). We further explore two key aspects: (1) The agreement of LLM-orders and semantic similarity orders, and (2) Points of maximum disagreement between these orders. To measure the ranking agreements, we calculate Kendall’s Tau correlation coefficient Kendall (1938) between the orders generated by ChatGPT, ChatGPT-Instruct, and Mistrial and orders sorted by semantic similarity scores, and the results are presented in fig. 6 and fig. 7. Surprisingly, for ChatGPT and ChatGPT-Instruct, we find that the two ranking orders have minimal monotonous relations with means approximating 00 and low standard deviations showing no agreement. For Mistral, we find a low agreement with a mean of $0.43$ score. These low and no agreements further verify that ranking by usefulness can be very different from ranking by semantic similarity. We also deep dive into cases with notable order variations to address (2). Section E.2 illustrates one such case in the "Guns" topic. We derive three observations. First, not all top-8 opinions by semantic similarity scores help predict the opinion. For example, the $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question. This is similar to the $18$-th opinion which is also less relevant. Meanwhile, several important opinions are deselected by the semantic-similarity-based method, such as the $6,3,4,10$-th ones, which are chosen by the LLM. The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership. Finally, by using LLMtop-$K$ order, the model predicts the opinion accurately, whereas the semantic similarity order leads to an incorrect prediction.
Opinions Order Analysis of LLMtop-$K$ Step.
@@
The performance difference of DIO-top8 and DIO-LLMtop8 in table 1 highlights that LLMs are sensitive to the chosen implicit personae opinions. An important question arises: Are LLMs also affected by the input order of implicit persona opinions in the ranking step (section 3)? Our discovery confirms sensitivity, but with reasonable overlap when $K$ is sufficiently large. We randomly select $300$ questions, shuffle implicit persona opinions four times with different seeds, and record four LLM ranking outputs for each. We also collect one more LLM ranking output by feeding implicit personae opinions in semantic similarity order. For each $K\in\{1,2,...,20\}$, we calculate the pairwise Overlap coefficient Vijaymeena and Kavitha (2016) among the five ranking outputs, averaging them as the LLM ranking consistency score for each $K$. The scores, shown in fig. 5, indicate that for $K\geq 8$, the ranking outputs overlap well with a score of .$\geq.6$ for both models. Despite this, is there substantial variance in model performance across random seeds? Our findings reveal no significant variance, with the variants statistically outperforming the baseline DIO-top8. Specifically, we assess ChatGPT and Mistral with DIO-LLMtop8 on $3$ out of $4$ random seeds, detailed in section B.6. The results demonstrate relatively small standard deviations in their performance, and critical values of $99\%$ CI of DIO-LLMtop8 under t-test for both models surpass DIO-top8, confirming that LLMtop8’s effectiveness is not due to randomness.
CoO versus CoT.
@@
Table 1 indicates that Chain-of-Thought (CoT) Kojima et al. (2022) slightly harms baseline DIO-top8 performance for ChatGPT and ChatGPT-Instruct. Conversely, our Chain-of-Opinion reasoning (CoO) enhances overall performance for all models. To investigate the consistency of CoT and CoO, we design an experiment with ChatGPT, DIO-top8 where we randomly select $100$ question-answer pairs and sample $5$ answers per pair using CoT and CoO, at $3$ different temperatures $0.3$, $0.6$, $0.9$. For each prompting technique, we measure the percentage of questions that all $5$ answers sampled have the same result, as the consistency score. The results are illustrated in fig. 2 showing that CoO brings slightly better consistent answers compared to CoT, especially when the temperature is high which verifies CoO potentially enhances the reliability of LLMs.
Dynamic Numbers of Opinions Analysis.
@@
Table 4 illustrates our analysis answering two research questions: (1) How frequent can’t LLMs answer the question? and (2) How do LLMs perform when more opinions than $K=8$ are provided in ChOiRe? Our findings show that, firstly, with $8$ opinions, GPT-4 exhibits the highest percentage of unanswered questions, while Mistral answers all the questions. Secondly, increasing the number of opinions beyond $8$ reduces this percentage across models, confirming our hypothesis regarding the lack of implicit personae opinions when fixing $K=8$ in section 3. Lastly, while including more opinions could possibly harm the performance of models, our answer consistency strategy enables LLMs to achieve the best results across three different K values.
Model ChatGPT ChatGPT-Inst GPT-4 Mistral
% of ITA of DIO-LLMtop8 + FEA + CoO 0.61 1.32 9.71 0.00
DIO-LLMtop8 + FEA + CoO 51.90 53.01 59.02 54.21
% of ITA of DIO-LLMtop10 + FEA + CoO 0.12 1.01 5.44 0.00
DIO-LLMtop10 + FEA + CoO 51.55 52.74 58.88 53.88
% of ITA of DIO-LLMtop12 + FEA + CoO 0.00 0.66 3.12 0.00
DIO-LLMtop12 + FEA + CoO 51.60 52.31 59.11 52.96
ChOiRe 52.21 53.26 59.30 54.43
Table 4: Extra analysis on ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. ITA stands for "Impossible To Answer".
6.2 Error Analysis
FEA Misses Key Explicit Personae.
Despite showing promising results in removing unuseful explicit personae depicted in table 3, we observe that LLMs sometimes misselect relevant personae. One such example is the top-left of section E.6. We observe that in this case, our annotators can’t grade a high FEA satisfaction score because "Education" and "Age" are also two important personae as they can influence one’s understanding of workplace dynamics significantly, which are deselected by ChatGPT.
LLMtop-$K$ Opinions Include Less Relevant Ones.
While LLMs generally demonstrate a commendable ability to rank implicit opinions by usefulness, as exemplified in section E.2, we also observe they frequently include less relevant, or even irrelevant opinions to the ranked list such as in section E.6-bottom. We attribute this to the challenge of this task, even for humans it might require substantial cognitive effort.
LLMs May Not Follow the Instructions.
Although ChatGPT and ChatGPT-Instruct demonstrate a robust ability to adhere to our instructions for opinion prediction via CoO, the same level of proficiency is not observed in Mistral and GPT-4, as shown in section E.6-top-right. We posit this disparity arises from the fact that ChatGPT and ChatGPT-Instruct excel in comprehending and executing human instructions, while GPT-4 excels primarily in generating coherent text.

### v5

@@ 4.1 COO in Prompting
$\rightarrow\hat{a}:=G_{\mathcal{M}}([E_{rel},I_{rel},Q])$ (4)
Eq. 3 is achieved in a chain of thoughts:
$(I_{1})$ Analyze the user’s demographics and ideology one by one to infer their social and environmental values.
$(I_{2})$
Analyze the user’s historical opinions one by one to infer their beliefs and norms from their social and environmental values.
$(I_{3})$
Which opinion is the user likely to choose?
Fig. 1 shows an example of VBN reasoning. Note that the instruction $(I_{1})$/$(I_{2})$ is skipped if $E_{rel}$/$I_{rel}$ is unavailable. Our human evaluation in Section 5.2 confirms LLMs are capable of reasoning about human values, beliefs, and norms. This reasoning offers two notable advantages. First, for each question, we ensure that the model explains and analyzes the provided personae one by one thoroughly without omitting any, resulting in more accurate predictions. Second, this method helps the model to output more consistent reasoning explanations, enhancing model reliability (Section 6.2).
Step 4. Answer consistency with dynamic numbers of opinions.
By fixing $K$ in the $LLMtop-K$ step, we observe that models such as GPT-4 OpenAI (2023b) may refuse to answer the question (Tab. 5). We attribute this to insufficient implicit personae provided. Inspired by Self-Consistency (SC) Wang et al. (2023b), our approach involves sampling multiple answers using different $K$ values for a given question. The most frequent answer, along with the explanation of the first correct answer, becomes the final prediction. Our method is distinct from SC which samples multiple answers with a fixed prompt. We only use three values of K $\{8,10,12\}$ for efficiency.
4.2 COO in Fine-tuning
During fine-tuning, we adopt COO’s Steps 1, 2, 3 in Section 4.1 to create COO data. Each sample is denoted as {$E_{rel},I_{rel},\text{Env. Values},\text{Beliefs and Norms},Q,a$} where the model learns to predict $a$ given the rest variables. Since these steps require LLMs with strong instruction-following capabilities, we use ChatGPT as the COO data processor.
5 Experiment
5.1 Experimental Settings
Dataset.
We experiment on OpinionQA dataset Santurkar et al. (2023) — the only open-sourced opinion QA dataset to date consisting of both user explicit and implicit personae designed for the assessment of alignment between LMMs’ opinions and human participants, encompassing a diverse range of $60$ US demographic groups. We also note the MRFP dataset Sun et al. (2023), which, due to privacy concerns, is unavailable for public access.
Dataset preprocessing.
Due to limited resources, we follow Hwang et al. (2023) to sample a subset of OpinionQA for evaluation. We randomly select $25$ users per topic for our experiments. We use $20\%$ of the implicit questions for each user as the implicit persona. For the remaining $80\%$ implicit questions, we randomly select a maximum of $15$ implicit questions for testing. Our sampling method results in a total of $375$ users and $5,603$ implicit evaluation question–answer pairs. Our subset is highly representative because we gather users from every topic and rigorous statistical tests further validate the significance of our results.
Baselines.
For prompting experiments, we use both closed- and open-source LLMs: ChatGPT OpenAI (2022), ChatGPT-it OpenAI (2023a), GPT-4 OpenAI (2023b), and Mistral-7B-it-v.02 Jiang et al. (2023) where each model performs all the COO’s steps. We compare COO with $5$ prompting methods: W/o persona, where LLMs are evaluated without user historical opinions, ideology, or demographic data; Demographic + Ideology + top8 Opinions (DIO-top8), introduced by Hwang et al. (2023) which achieves state-of-the-art results on OpinionQA at that time; DIO-top8 + CoT is the Chain-of-Thought (CoT) prompting Kojima et al. (2022) version of DIO-top8 by appending "answer the following question step-by-step" to prompts; DIO-top8 + SC is when we apply the Self-Consistency Wang et al. (2023b) with CoT to DIO-top8; DIO-top8 + Self-refine Madaan et al. (2023) interactively feedbacks and refines the answers by LLMs. For GPT-4, we only run the main experiment and use ChatGPT for FEA and LLMtop-$K$ steps due to our limited budget. We provide all the prompt templates and cost analysis in Appendix C, implementation details in Appendix A, and more baselines in Section B.2.
For fine-tuning, we use ChatGPT to perform COO’s steps 1, 2 ($K=8$), 3 as noted in Section 4.2 on a training set of $30,000$ samples randomly selected from OpinionQA which are different from our $5,603$ test ones. We then fine-tune and evaluate GPT-2 models (base, large) Radford et al. (2019) and FlanT5 models (base, large) Chung et al. (2024). The details are shown in Appendix A.
Prompting Fine-tuning
Model ChatGPT ChatGPT-it Mistral-7B-it-v0.2 GPT-2-base GPT-2-large FlanT5-base FlanT5-large
W/o personae 46.60 / 65.72 44.91 / 63.60 41.24 / 59.54 36.28 / 52.62 21.94 / 39.11 48.98 / 68.33 39.83 / 58.43
@@ 5.1 Experimental Settings
+ VBN (Step 3) 51.38 / 70.32 52.61 / 71.90 53.59 / 73.46 24.79 / 43.50 28.73 / 47.09 58.21 / 77.80 56.87 / 76.18
DIO-LLMtop8 (Step 2) 51.03 / 70.31 52.80 / 72.60 45.86 / 64.98 22.65 / 41.12 28.86 / 47.60 57.97 / 77.46 58.20 / 77.56
+ FEA (Step 1) 51.19 / 70.69 52.97 / 72.84 45.23 / 64.73 25.05 / 44.41 29.54 / 48.66 57.45 / 77.13 59.00 / 78.46
+ FEA + VBN 52.16 / 71.90 53.08 / 72.92 54.56 / 74.37 26.17 / 45.92 30.21 / 49.63 59.62 / 78.87 60.13 / 78.92
COO (ours) 52.66$\dagger$ / 72.75$\dagger$ 53.58$\dagger$ / 73.80$\dagger$ 54.40$\dagger$ / 74.26$\dagger$ 26.17$\dagger$ / 45.92$\dagger$ 30.21$\dagger$ / 49.63$\dagger$ 59.62$\dagger$ / 78.87$\dagger$ 60.13$\dagger$ / 78.92$\dagger$
% w/ best baseline + 4.11 / + 4.43 + 2.91 / + 2.68 + 2.37 / + 2.57 + 23.26 / + 18.84 + 21.13 / + 17.55 + 8.40 / + 5.18 + 9.45 / + 5.52
Table 2: Accuracy (Acc) / Collapsed Accuracy (CAcc) experimental results. FEA, LLMtop8, and VBN are COO’s Steps 1 (explicit), 2 (implicit), and 3 (reasoning). $\dagger$ denotes our method outperforms baselines with p-value < 0.01 under t-test (Tab. 10).
Automatic metrics.
We employ Accuracy (Acc) and Collapsed Accuracy (CAcc)222is a relaxed accuracy wherein the choices of MCQ questions ($\geq$ 4 choices) are collapsed to become 2 choices. as the evaluation metrics following Hwang et al. (2023). Note that Precision/Recall/F1 is not applicable in our task, since the numbers of answer choices are not the same for all the OpinionQA samples.
Human metrics.
Human evaluations are crucial due to the absence of automated metrics assessing LLMs’ performance in FEA, LLMtop-$K$, and VBN steps of COO. Therefore, we conduct our human assessments to address these research questions: (1) LLMs’ effectiveness in filtering unnecessary explicit personae; (2) LLMs’ proficiency in ranking implicit personae opinions; (3) LLMs’ ability to explain answers via VBN. We randomly select $100$ answers generated by COO with ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. We then hire $3$ excellent undergraduates who are native English speakers as annotators. For FEA and LLMtop-$K$ steps, each annotator is instructed to rate on a 1-3 scale (3 is best) via the Satisfaction criterion defined as how well the algorithm of LLMs performs in filtering/ranking, subjectively. To answer (3), we use two criteria named Reasonableness measuring how well the LLMs reason with the VBN explanations, and Follow the Instruction assessing the capability of LLMs in following our instruction to explain and predict the opinions. Three annotators are also guided to rate the criteria on a 1-3 scale. Each metric’s final score is the average of three annotators’ scores. The scoring instructions are in Section D.1 and the inter-annotators’ agreement is assessed by Krippendorff’s alpha Krippendorff (2011).
5.2 Main Results
We outline the prompting, fine-tuning, fine-grained, and human evaluation results of COO.
Prompting results.
Tab. 2 shows our main experimental outcomes. For GPT-4, it attains 57.98% Acc with DIO-top8 and 59.42% with COO, establishing a strong SOTA result surpassing the previous of 53.74% Hwang et al. (2023). Overall, COO delivers the best results with significant improvements of 2-4% Acc for all benchmarked LLMs with ChatGPT securing the most gain of 4.11%. Its component steps consistently enhance baselines with VBN > LLMtop8 > FEA on average.
Among prompting methods, we observe that naïve CoT helps Mistral slightly but harms ChatGPT and ChatGPT-it, while SC improves all. We attribute this to the inconsistency and unreliability of CoT reasoning (4), and the challenge of this task. Meanwhile, COO’s VBN reasoning consistently improves, verifying the effectiveness of requiring LLMs to explicitly analyze all the personae through values, beliefs, and norms. Additionally, self-refine consistently lowers model performance, indicating that multiple refinement rounds may be counterproductive for this complex task as these rounds may amplify the model’s inherent biases Xu et al. (2024b), leading to more biased predictions.
Among models, notably, for GPT-4, we use ChatGPT for FEA and LLMtop-$K$ steps, showcasing the strength of a weaker model that enhances a stronger one. Finally, ChatGPT, ChatGPT-it, and Mistral show improvements by selecting only $4.79/12$ and $5.59/12$, $8.83/12$ explicit personae on average. This suggests that over half of explicit personae are noisy for model opinion prediction.
Fine-tuning results.
Tab. 2 presents our fine-tuning results. Overall, leveraging COO’s FEA, LLMtop-$K$, and VBN extra variables results in significant improvements for both decoder-only and encoder-decoder models, with average gains of 22.20% for GPT-2s and 8.93% for FlanT5s. Among COO’s steps, FEA contributes the least, while LLMtop-$K$ and VBN extra variables drive more substantial gains across most models.
The VBN extra variables can be seen as distilled knowledge from ChatGPT, intuitively enhancing fine-tuning results. However, notably, COO’s FEA and LLMtop-$K$, which focus on selecting relevant explicit and implicit personae, already deliver substantial improvements across all models, bringing FlanT5-large’s performance on par with GPT-4. This verifies our hypothesis that removing irrelevant personae is necessary for high performance.
Finally, GPT-2-base performs surprisingly well even without user demographics and ideology, possibly due to potential contamination Sainz et al. (2023) with public polling data from OpinionQA.
Fine-grained Results.
We compare COO with the DIO-top8 baseline across 15 OpinionQA topics to assess its performance in detail. Overall, we see COO consistently outperforms DIO-top8 in most topics, with the largest improvements in “View on gender” (+17.69% with Mistral), “Autonomous vehicles” (+13.49 with GPT-2), and “Misinformation” (+11.61 with ChatGPT). These gains further emphasize COO’s effectiveness in enhancing LM performance on social and belief-driven topics, especially those involving complex reasoning. Full results are provided in Appx. Tab. 6.
Human evaluation results.
Model FEA Satis. LLMtopK Satis. VBN Rea. VBN FI
ChatGPT $2.56_{\alpha=0.74}$ $\textbf{2.32}_{\alpha=0.68}$ $2.85_{\alpha=0.88}$ $\textbf{2.91}_{\alpha=0.90}$
ChatGPT-it $\textbf{2.64}_{\alpha=0.71}$ $2.28_{\alpha=0.65}$ $2.87_{\alpha=0.90}$ $\textbf{2.95}_{\alpha=0.87}$
GPT-4 $\textbf{2.90}_{\alpha=0.91}$ $2.21_{\alpha=0.77}$
Mistral $2.31_{\alpha=0.65}$ $2.12_{\alpha=0.64}$ $2.58_{\alpha=0.68}$ $2.16_{\alpha=0.55}$
Table 3: Human evaluation results. LLMs perform adequately in COO’s FEA step, and excel in VBN reasoning, but face challenges with the LLMtop-K step. $\alpha$ denotes the Krippendorff’s alpha.
@@ 5.2 Main Results
For COO’s FEA and LLM ranking steps, from Tab. 3, ChatGPT and ChatGPT-it generally achieve similar performance and are better than Mistral: ChatGPT excels slightly in ranking while ChatGPT-it performs slightly better in performing FEA. Three models are proficient in FEA but struggle with the ranking task where the common error is misplacing relevant opinions, due to this task’s complexity. Second, four models effectively generate VBN reasoning thoughts, and GPT-4 performs the best. Finally, ChatGPT and ChatGPT-it follow our instructions to explain and analyze the explicit and implicit personae provided one by one with VBN significantly better than GPT-4 and Mistral, achieving nearly perfect scores of $3$. Our hypothesis is they are optimized for following instructions, while GPT-4 is optimized for completing texts.
@@ 6 Discussion
We discuss the main analyses here, including COO generalization (Section 6.1) and COO’s Steps sequentially. Extra analyses are supplemented in Appendix B.
Method ChatGPT
W/o personae 46.60
COO w/o personae (Step 4 activated) 47.79
DIO-top8 w/o explicit 49.22
COO w/o explicit (Steps 2, 3 (PBN), 4 activated) 51.66
DIO-top8 w/o implicit 47.16
COO w/o implicit (Steps 1, 3 (EV), 4 activated) 50.13
Table 4: COO’s results with missed persona(e) with ChatGPT. In all scenarios, COO outperforms the baselines significantly.
6.1 COO in Missing Personae Scenarios
COO demonstrates strong generalizability even when explicit, implicit, or both types of personae are missing. Specifically, in the absence of both, COO reduces to Self-consistency Wang et al. (2023b); without explicit personae, only Step 1 of COO is skipped; without implicit personae, Step 2 is omitted. As shown in Tab. 4, COO consistently outperforms the leading baseline by an absolute margin of 2-3% in all scenarios. Finally, its steps can be generalizable to other tasks, see Section B.8.
6.2 Method Analysis
FEA: ablation study.
@@ 6.2 Method Analysis
To gauge the impact of removing irrelevant explicit personae (FEA), we experiment with applying FEA exclusively to the baseline DIO-top8 Hwang et al. (2023), denoted as DIO-top8 + FEA in Tab. 2. We observe a 1-2% Acc performance boost on ChatGPT, ChatGPT-it, and Mistral respectively. This underscores the effectiveness of eliminating irrelevant explicit personae in improving the model prediction.
FEA: irrelevant personae distribution.
@@
To understand the explicit personae filtered by LLMs across various topics, we document the top 3 removed personae in Section B.3. "Citizenship" is observed to be the most frequently removed attribute, followed by "Race". This could be due to LLMs treating these as sensitive information, prioritizing respect and unbiased text generation. Another explanation may be the lack of correlation between citizenship/race and opinions in the OpinionQA dataset. Additionally, we also see that ChatGPT often categorizes “Marital status" as non-useful, ChatGPT-it commonly removes “Frequency of religious attendance", and “Gender" got removed by Mistral, revealing potential biases in LLMs.
LLMtop-$K$: compared to semantic top-$K$.
From Tab. 2, DIO-LLMtop8 outperforms DIO-top8 by $1-4\%$ accuracy on ChatGPT, ChatGPT-it, and Mistral, confirming that prioritizing usefulness over semantic similarity improves model prediction. To further understand the difference between semantic similarity orders and usefulness orders, we discuss (1) the agreement of LLM orders and semantic similarity orders, and (2) maximum disagreement points between these orders.
In tackling (1), we calculate Kendall’s Tau coefficient Kendall (1938) between the orders generated by ChatGPT, ChatGPT-it, Mistrial, and semantic similarity orders, and the results are presented in Appx. Fig. 4. Surprisingly, for ChatGPT and ChatGPT-it, we find that the two ranking orders have no agreement with means approximating 00. For Mistral, we observe a low agreement with a mean of $0.43$ score. These low and no agreements further verify that ranking by usefulness can be very different from ranking by semantic similarity.
@@
For answering (2), Section E.2 illustrates one such case in the "Guns" topic. We observe that not all top-8 opinions by semantic similarity scores help predict the opinion. For example, the $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question. This is similar to the $18$-th opinion which is also less relevant. Meanwhile, several important opinions are deselected by the semantic-similarity-based method, such as the $6,3,4,10$-th ones, which are chosen by the LLM. The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership. Finally, by using the LLMtop-$K$ order, the model predicts the opinion accurately, whereas the semantic similarity order leads to an incorrect prediction.
LLMtop-$K$: the order of input opinions.
We study the (1) sensitivity and the (2) performance variance of LLMs to the order of input implicit personae in the LLMtop-$K$ step.
@@
To address (1), our discovery confirms sensitivity, but with reasonable overlap when $K$ is sufficiently large ($K\geq 8$). We randomly select $300$ questions, shuffle implicit persona opinions four times with different seeds, and record four LLM ranking outputs for each. We also collect one more LLM ranking output by feeding implicit personae opinions in semantic similarity order. For each $K\in\{1,2,...,20\}$, we calculate the pairwise Overlap Coefficient (OC) Vijaymeena and Kavitha (2016) among the five ranking outputs, averaging them as the LLM ranking consistency score for each $K$. The scores, shown in Appx. Fig. 3, indicate that for $K\geq 8$, the ranking outputs overlap well with a score of .$\geq.6$ for both models.
For (2), we find no significant performance variance. Specifically, we assess ChatGPT and Mistral with DIO-LLMtop8 on $3$ out of $4$ random seeds, detailed in Section B.5. The results demonstrate relatively small standard deviations in their performance, and critical values of $99\%$ CI of DIO-LLMtop8 under t-test for both models surpass DIO-top8, confirming that LLMtop8’s effectiveness is not due to randomness.
VBN: compared to CoT.
@@
Tab. 2 indicates that Chain-of-Thought (CoT) Kojima et al. (2022) slightly harms the performance for ChatGPT and ChatGPT-it. Conversely, our Value-Belief-Norm (VBN) reasoning enhances performance for all models. To investigate the consistency of CoT and VBN, we design an experiment with ChatGPT, DIO-top8 where we randomly select $100$ question-answer pairs and sample $5$ answers per pair using CoT and VBN, at $3$ different temperatures $0.3$, $0.6$, $0.9$. We measure the percentage of questions that all $5$ answers sampled have the same result, as the consistency score. The results are illustrated in Appx. Fig. 2 showing that VBN brings better consistent answers compared to CoT, especially when the temperature is high verifying VBN potentially enhances the reliability of LLMs.
Answer consistency with dynamic opinions.
@@
We study (1) how frequently LLMs are unable to answer the question and (2) the impact on performance when more than $K=8$ opinions are provided. Tab. 5 provides the results. We find that with $8$ opinions, GPT-4 exhibits the highest percentage of unanswered questions, while Mistral answers all the questions. Increasing the #opinions beyond $8$ reduces this percentage across models, confirming our hypothesis regarding the lack of implicit personae opinions when fixing $K=8$ in Section 4. Lastly, while increasing $K$ could harm the model performance, COO’s answer consistency enables LLMs to achieve the best results across K values.
Model ChatGPT ChatGPT-it GPT-4 Mistral
ITA of DIO-LLMtop8 + FEA + VBN 0.20 0.91 3.40 0.00
DIO-LLMtop8 + FEA + VBN 52.16 53.08 59.11 54.56
ITA of DIO-LLMtop10 + FEA + VBN 0.00 0.00 1.43 0.00
DIO-LLMtop10 + FEA + VBN 51.89 52.90 58.88 53.62
ITA of DIO-LLMtop12 + FEA + VBN 0.00 0.00 0.00 0.00
DIO-LLMtop12 + FEA + VBN 51.60 52.03 59.18 54.21
COO 52.66 53.58 59.42 54.40
Table 5: Percentage of “Impossible To Answer” (ITA) (%) observed with corresponding performance during generation.